OpenStudy (anonymous):

let sinA =3/5 in QII and sinB = -5/13 in QIII find sin(A-B), cos(A-B),and tan (A-B). What quadrant does (A-B) terminate in?

OpenStudy (john_es):

\[\sin(A-B)=\sin A\cos B-\cos B\sin A=-3/5\cdot\sqrt{1-(5/13)^2}+\sqrt{1-(3/5)^2}\cdot5/13\] \[\cos(A-B)=\cos A\cos B+\sin B\sin A=\sqrt{1-(3/5)^2}\cdot\sqrt{1-(5/13)^2}-5/13 \cdot 3/5\] \[\tan(A-B)=\sin(A-B)/\cos(A-B)\]

OpenStudy (john_es):

The last fraction of sin(A-B) is 5/13.

OpenStudy (john_es):

The final angle is in the Q IV, because the sin(A-B)<0 and cos(A-B)>0.